[figure: 9.7 cm]

HIGH COERCIVITY OF Nd-Fe-B THIN FILMS PREPARED BY RF-SPUTTERING WITH SUBSEQUENT HEAT-TREATMENT

L.K.E.B. Serrona, T. Okuda, R. Fujisaki, A. Sugimura, N.Adachi, I. Sakamoto, H. Ohsato, A. Nakanishi^{a}, and M. Motokawa^{b} Nagoya Institute of Technology, Nagoya, 466-8555 Japan a Sumitomo Special Metals Co., Osaka, 618-0013 Japan b Institute for Materials Research, Tohoku University, Sendai, 980-8577 Japan

We prepared Nd − F e − B thin films by RF-sputtering with subsequent heat-treatment. The films possess perpendicular magnetic anisotropy, H_{c} ≥ 15kOe, (BH)_{max} ≥ 20MGOe, M_{r}/M_{s} ≠ 1. The initial magnetization curve suggests that the coercivity mechnism in these films is nucleation type. The as-deposited films are in amorphous state and show soft magnetic properties with strong in-plane and weak out-of-plane anisotropy. In these films heat-treated at 650^{∘}C for 30min, Nd_{2}F e_{14}B phase crystallizes with c-axis orientation along the film normal. SEM observation revealed that the as-deosited films have dense, continous structure, on the other hand, the heat-treated films have the columnarlike texture consisting of many fine crystallites. The fine structure that originates weak out-of-plane anisotropy in as-deosited film may concern with the fine structure which that results in excellent, hard, magnetic characteristics in the heat-treated films.

[figure: 13.4 cm]

Subject category :

2. Magnetic Films, Surfaces and Multilayers

Presentation mode : oral

Corresponding author : T. Okuda

Address for correspondence : Nagoya Institute of Technology Gokiso-cho, Showa-ku, Nagoya 466-8555 Japan

Email address : okudatks@mse.nitech.ac.jp